OpenStudy (anonymous):

Math masters help! :D Simplify the rational expression. State any restrictions on the variable. n^4 -11n^2 + 30 / n^4 -7n^2 + 10

OpenStudy (anonymous):

well i know the n^4s cancel eachother out

OpenStudy (apoorvk):

No they dont cancel out each other, @emmy

OpenStudy (jlastino):

Hmmm n= +- sqrt 6

OpenStudy (anonymous):

@emmy82650 thats incorrect

OpenStudy (apoorvk):

Factorise the numerator and denominator by assuming n^2=t, and then like a quadratic, math4life, and then replace t=n^2

OpenStudy (anonymous):

okay so \[n ^{4}-11n ^{2}+30 \over n ^{4}-7n ^{2}+10\]

OpenStudy (anonymous):

that's all i've got then lol

jhonyy9 (jhonyy9):

(n^2 -5)(n^2 -6) ---------------- = (n^2 -6)/n^2 -2) (^2 -5)(n^2 -2)

jhonyy9 (jhonyy9):

so can you continue it ?

OpenStudy (anonymous):

@apoorvk \[n^2-6/n^2-2\] im pretty sure this is the equation you get

OpenStudy (anonymous):

@jhonyy9 lol i just got that equation

OpenStudy (anonymous):

@zarkon can you help solve this?

OpenStudy (amistre64):

the restrictions are when the denominator goes zero

jhonyy9 (jhonyy9):

first you need push these restrictions - do you know it ?

OpenStudy (amistre64):

n^4 -7n^2 + 10 = 0 ; when n^2 = 5,2 or when n = sqrt(5), -sqrt(5), sqrt(2), -sqrt(2)

OpenStudy (anonymous):

@johnyy9 @amistre64 \[n \neq 5 \] \[n \neq -2\] ?

OpenStudy (amistre64):

not n, n^2

OpenStudy (amistre64):

you can clean up the ns by saying x=n^2, x^2=n^4 x^2-7x+10 = 0 x = ..... = n^2

jhonyy9 (jhonyy9):

n^2 not can being equal 2

OpenStudy (anonymous):

@amistre64 would this be my answer? \[n^2-6/n^2 -2; n \neq 5, n \neq 2\]

OpenStudy (amistre64):

no

OpenStudy (amistre64):

n^2 - 7 +10/n^2 maybe

OpenStudy (amistre64):

but that just makes things wierd; sub in x (or your favorite algebraaic symbol) for n^2 and solve ....

OpenStudy (anonymous):

@amistre64 n^4 - 11n² + 30 / n^4 - 7n² + 10...........Let n² = a a² - 11a + 30 / a² - 7a + 10 = (a - 6)(a - 5) / (a - 2)(a - 5) = (a - 6) / (a - 2) = (n² - 6) / (n² - 2) Restriction : n² ≠ 6 and n² ≠ 2

jhonyy9 (jhonyy9):

so just n^2 not can be equal 2 than n not can being equal +/- sqrt2

jhonyy9 (jhonyy9):

these are those restrictions for